OpenStudy (anonymous):

12 + 0.35x = 20.05 A. 91.5 B. 57.3 C. 2.8175 D. 23

sam (.sam.):

12+0.35x=20.05 Since 12 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 12 from both sides. 0.35x=-12+20.05 Add 20.05 to -12 to get 8.05. 0.35x=8.05 Divide each term in the equation by 0.35. (0.35x)/(0.35)=(8.05)/(0.35) Cancel the common factor of 0.35. x=(8.05)/(0.35) Divide 8.05 by 0.35 to get 23. x=23

OpenStudy (anonymous):

You're awesome .

sam (.sam.):

thank you :)